OpenStudy (anonymous):

The population f(x), in millions, of State A of a country after x years is represented by the function shown below:

OpenStudy (anonymous):

\[f(x)=2(1.08)^{t}\] The graph shows the population g(x), in millions, of State B of the country after x years:

OpenStudy (anonymous):

Which conclusion is correct about the population of State A and State B? (A) The original population of State B was half of the original population of State A. (B) The original population of State A was half of the original population of State B. (C) The original population of State A was equal to the original population of State B. (D) The original population of State B was one-fourth of the original population of State A.

OpenStudy (anonymous):

@cwrw238 @hartnn @Zarkon @iambatman @ganeshie8 I need help with this. Please explain this problem to me.

OpenStudy (anonymous):

I think you have to graph the function of State A and then compare the graphs.

OpenStudy (anonymous):

Doing in my head, the function of f(x) = g(x)'s graph making letter C the correct answer.

OpenStudy (anonymous):

I don't think C could be the answer.

OpenStudy (anonymous):

\[f(x) = 2(1.08)^{t}\] Say 0 years has passed, t = 0 \[f(x) = 2(1.08)^{0}= 2(1)=2\]

OpenStudy (anonymous):

f(x) graph, (x, y) = 0, 2 g(x) graph (x, y) = 0, 2 Just to be sure, Imma get somebody @Directrix

OpenStudy (anonymous):

\[f(x)=2(1.08)^{t}\]

OpenStudy (anonymous):

is equal to \[f(x)= 2(1)^{0.08t}\]

OpenStudy (anonymous):

@iGreen @paki I really need help on this.

OpenStudy (anonymous):

@pooja195 can u help me plz?

pooja195 (pooja195):

@hartnn

hartnn (hartnn):

yes original population of both states = 2 they are equal :)

OpenStudy (anonymous):

Can you explain? I dont really understand it tho. @Sadworld said that it was C also.

hartnn (hartnn):

original population means population at t=0 for sate A , when t = 0 \(2 (1.08)^0 = 2\) because anything^0 =1 (except 0)

hartnn (hartnn):

for state B, its the y intercept, the point at which graph cuts y axis which is clearly 2 :)

OpenStudy (anonymous):

Ohhhh, ok. Thank you so much @Sadworld @hartnn @pooja195!! :)

hartnn (hartnn):

thanks @pooja195 :)

pooja195 (pooja195):

lol i didnt do anything but your welcome

pooja195 (pooja195):

@hartnn yr welcome :)